Find a polynomial of degree 5 and zeros of –6, –2, 0, 2, and 6
every time i try, i get x^5+12x^3+12x^2-144x

Answer by ewatrrr(24785)

 
Hi,
degree 5 and zeros of –6, –2, 0, 2, and 6
x(x -6)(x+6)(x-2)(x+2)= x(x^2-36)(x^2-4) = (x^3 - 36x)(x^2 - 4)